Calculate the pH and pOH of the solutions with the following hydronium ion [H3O+] or hydroxide ion [OH–] concentrations. Determine which solutions are acidic, basic or neutral.
(1.a) [OH–] = 8.2 × 10–11 M
(1.b) [OH–] = 7.7 × 10–6 M
(1.c) [H3O+] = 3.2 × 10–4 M
(1.d) [H3O+] = 1.0 × 10–7 M
Answer #1

(a) [OH-] = 8.2 X10 M pOH = -log.com-J = -log(8.2x10) = 10.086 ÞH = 14 - poh - 14- 10.086 = 3.914 [got] or [HT] = 10PH = 103

(1c) [H₃O+ ] = 3,2 x 100 M pH = -log[H₃O+] = 3.49 pOH = 14- pH = 10.51 TOH-] = 10 Pot = 10 10.51 - 3.09 x 100 m The solution

Note :- pH = - Da Cof Сн,oto Сон-] = хоч рн + Бон = јч рон = - log сон-1 Сңot] = БРң Сон:] - - Бон pH > 7 pH <7 Base Aid